OpenStudy (anonymous):

Use the change base formula to evaluate log4 27. The convert log4 27. Then convert log4 27 to a logarithm in base 2. Round to the nearest thousandth. A.2.377; log2 13.5 B.1.431; log2 13.5 C.3.296; log2 5.196 D.2.377: log2 5.196 I have already did most of the math and came up with A or D. I just can't figure out the second part of the question

OpenStudy (anonymous):

I hate this math topic lol

OpenStudy (anonymous):

its better than geometry

OpenStudy (anonymous):

tru dat tru dat

OpenStudy (cwrw238):

change of base formula log (b) x = log(d) x / log(d) b

OpenStudy (anonymous):

yeah i got that log 27/ log 4 = 2.377

OpenStudy (anonymous):

I cant figure out the second part of the answer.

Nnesha (nnesha):

i guess they mean \[\log _{2} 27\]

OpenStudy (cwrw238):

i'm confused by this question i must admit i get log4 27 equal to (1/2) log2 27 or log2 27^(1/2)

OpenStudy (anonymous):

I don't know what they mean? @Nnesha. and where are you getting (1/2) from? @cwrw238

OpenStudy (cwrw238):

log4 27 = log2 27 / log2 4 by the change of base formula and since log2 4 = 2 log4 27 = (½) log2 27

OpenStudy (cwrw238):

thats a half log2 27

OpenStudy (cwrw238):

= log2 27^(1.2) = log2 5.196

OpenStudy (cwrw238):

so its D

OpenStudy (anonymous):

where did 1.2 come from?

OpenStudy (cwrw238):

sorry thats 1/2

OpenStudy (cwrw238):

27^(½) = sqrt27 = 5.196

OpenStudy (cwrw238):

i used one of the rules of logs n log a = log a^n

OpenStudy (anonymous):

whats a and n?

OpenStudy (cwrw238):

thats a general formula a and n stand for any numbers if you put n = ½ and a = 27

OpenStudy (cwrw238):

you get the above result

OpenStudy (anonymous):

well no because you put n as a question mark

OpenStudy (cwrw238):

thats open study problem it was meant to be a half

OpenStudy (anonymous):

yea but 27^? doesn't =1/2

OpenStudy (cwrw238):

27 to the power half = square root of 27 = 5.196

OpenStudy (anonymous):

okay the root sign does. i see now.

OpenStudy (cwrw238):

good

OpenStudy (anonymous):

Thank you :)

OpenStudy (cwrw238):

yw